A particular fruit's weights are normally distributed, with a mean of 536 grams and a standard deviation of 17 grams.

If you pick one fruit at random, what is the probability that it will weigh between 482 grams and 519 grams
Answer #1

z score for 482 = (482 - 536)/17 = -3.18

z score for 519 = (519 - 536)/17 = -1

Hence,

P(482 < X < 519)

= P(-3.18 < z < -1)

= 0.1579